Consider the following data for compound A:

(delta)Hf(kj/mol) S(J/K*mol)

A(s) -224.57 32.36

A(l) -172.45 109.55

Use these data to determine the normal freezing point of compound A. Report your answer to 3 significant figures and degrees Celsius)

Freezing point =
Answer #1

Change in enthalpy = -172.45 -(-224.57) = 52.12 kJ/mol

Change in entropy = 109.55 - 32.36 = 77.19 J/mol K

at phase equilibrium delta G = 0

delta G = delta H - T x delta S

0 =   52.12 kJ/mol - T* 77.19 J/mol K

T = 52120 / 77.19

= 675.21 K

= 402 oC